# All for One, and One for All: UrbanSyn Dataset, the third Musketeer of Synthetic Driving Scenes

### v1

@@ V Conclusions
In this paper, we introduce UrbanSyn, a diverse, compact, and photorealistic dataset of synthetic images capturing urban driving scenes. It comes with pixel-wise ground truth for depth estimation, and semantic and instance segmentation. A 2D BB and a degree of occlusion complement instance labels. UrbanSyn brings differences from GTAV and Synscapes in terms of road layouts, background content, and traffic participants. All of them form the Three Musketeers dataset. We have demonstrated the value of UrbanSyn individually and as part of the Musketeers by establishing new benchmarks on synth-to-real UDA for onboard semantic segmentation, where we have used state-of-the-art methods with core differences (HRDA, co-training) and a variety of real-world target domains (Cityscapes, Mapillary Vistas, BDD100K). We make UrbanSyn openly and freely accessible to the research community. Therefore, we expect that the research community can use UrbanSyn not only for tasks related to semantic segmentation, but also for instance segmentation, object detection, and monocular depth estimation. We plan to develop new pixel-level self-labeling protocols taking advantage of the class and instance labels provided by UrbanSyn.

### v2

@@ V Conclusions
This paper introduces UrbanSyn, a diverse, compact, and photorealistic dataset of synthetic images capturing urban driving scenes. It comes with pixel-wise ground truth for depth estimation and semantic and instance segmentation. A 2D BB and a degree of occlusion complement instance labels. UrbanSyn brings differences from GTAV and Synscapes regarding road layouts, background content, and traffic participants. All of them form the Three Musketeers dataset. We have demonstrated the value of UrbanSyn individually and as part of the Musketeers by establishing new benchmarks on synth-to-real UDA for onboard semantic segmentation, where we have used state-of-the-art methods with core differences (HRDA, co-training) and a variety of real-world target domains (Cityscapes, Mapillary Vistas, BDD100K). We make UrbanSyn openly and freely accessible to the research community. Therefore, we expect that the research community can use UrbanSyn not only for tasks related to semantic segmentation but also for instance segmentation, object detection, and monocular depth estimation. As future work, we plan to work on synth-to-real UDA for the instance segmentation task, where pixels must not only be assigned a semantic class but also an individual identifier (e.g., to distinguish one vehicle from another). For this task, it is essential that UrbanSyn not only provides class labels but instance labels too. In this regard, UrbanSyn offers fine-grained instance information as distinguishing instances of motorbikes/bikes from their human riders in contrast to Synscapes’ instances (where the rider and the motorbike/bike form the same instance) and GTAV-based instances (no bikes/cyclists are included, and motorbike riders are treated as pedestrians). In addition, as future work, we want to explore how UrbanSyn can help in synth-to-real active learning procedures, where, in contrast to UDA, this approach allows for a certain budget to label a few real-world training images by hand. The challenge lies in selecting which images to allocate that budget.
Accessing the dataset
In order to download UrbanSyn, users have two options, either accessing the UrbanSyn web page, https://urbansyn.org, or accessing the popular HuggingFace services, https://huggingface.co/datasets/UrbanSyn/UrbanSyn. Both sites include all the information required to download the data, a description of their content, and visual examples.
Acknowledgements
This work has been supported by the Spanish grants Ref. PID2020-115734RB-C21 (ADA/SSL-ADA subproject) and PID2020-115734RB-C22 (ADA/PGAS-ADA subproject), both funded by MCIN/AEI/10.13039/501100011033. Antonio M. López acknowledges the financial support to his general research activities given by ICREA under the ICREA Academia Program. CVC’s authors acknowledge the support of the Generalitat de Catalunya CERCA Program and its ACCIO agency to CVC’s general activities. Jose A. Iglesias-Guitian acknowledges the financial support to his general research activities given by UDC-Inditex InTalent programme, the Spanish Ministry of Science and Innovation (AEI/RYC2018-025385-I), and Xunta de Galicia (ED431F 2021/11, EU-FEDER ED431G 2019/01).